find an equation for a line that goes through(-3,-1) and has slope 6
Answer by nerdybill(7384)

find an equation for a line that goes through(-3,-1) and has slope 6
.
Plug into the "point-slope" form:
y - y1 = m(x-x1)
y - (-1) = 6(x-(-3))
y + 1 = 6(x+3)
y + 1 = 6x+18
y = 6x + 17 (this is what they're looking for)